AnakKrakatau1883:

Simplify 4 x times the square root of quantity 3 x end quantity minus x times the square root of quantity 3 x end quantity minus 2 x times the square root of quantity 3 x end quantity period. X times the square root of quantity 3 x end quantity x times the square root of quantity 9 x end quantity 2 x times the square root of quantity 6 x end quantity 2 x times the square root of quantity 6 x cubed

Spectrum:

what is each quantity equal to

jhonyy9:

do you think on this in this way ? \[4x \sqrt{3x} - x \sqrt{3x} - 2x \sqrt{3x} = ?\]